# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

### v5

@@ Abstract
Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures. However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations , the two major ones being: i) the definition of the QM and MM regions as well as the description of their coupling, and ii) the choice of the methods and levels of approximation made to describe each region
This paper addresses the second point and develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation. We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble. We start by recalling the expression of the adiabatic equilibrium density matrix for a mixed system made of $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles. Then, we propose a variational formulation of the Helmholtz free energy in terms of the full , non-equilibrium, QM/MM density matrix. Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities. Therefore, this paper generalizes both cDFT and electronic DFT (eDFT) to QM/MM systems.
@@ I Introduction
Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications. Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion due to the typically large gaps in electronic energies. In this context, because of the large mass of nuclei with respect to that of electrons, it is physically relevant to maintain electrons at the quantum level and to treat nuclei as classical particles. Nevertheless, the forces acting on the nuclei are computed from the solutions of the quantum electronic ground state, i.e. from the solutions of the Born Oppenheimer Hamiltonian. This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999) which recovers the classical nuclear entropic contributions through either classical Monte Carlo algorithms or classical trajectories under the ergotic hypothesis.
@@
Nevertheless, it is important to keep in mind that nuclear quantum effects play a significant role for light nuclei and can substantially alter the rates of certain chemical processes, such as proton transfer reactionsBała et al. (1996); Hanna and Kapral (2005); Kim and Hammes-Schiffer (2006); Costentin, Robert, and Savéant (2010); Hammes-Schiffer and Stuchebrukhov (2010). Moreover, such proton transfers can also be strongly influenced by the solvation effects, specially when polar solvents are usedKapral (2015). In this context, two fundamentally different yet interconnected nuclear quantum effects are at work: (i) the intrinsic delocalization of nuclear wave packets due to the quantum nature of nuclei, and (ii) the nonadiabatic couplings that arise between various adiabatic electronic states. Nuclear delocalization can be captured through explicit nuclear wave functions (see, for example, Ref. Beck et al., 2000 for a review of time-dependent mean-field approaches, and Ref. Hammes-Schiffer, 2021 for the nuclear-electron orbital approach) or via path-integral formalisms such as ring-polymer molecular dynamicsHabershon et al. (2013). Nonadiabatic couplings, on the other hand, are commonly addressed using surface hopping methods based on classical nuclear trajectoriesTully (1990); Hammes-Schiffer and Tully (1994); Ben-Nun and Martı’nez (2000); Li et al. (2005); Richardson and Thoss (2013); Ananth (2013), modified time-dependent mean-field techniquesWorth et al. (2008), semiclassical expansions of the density matrix and its Wigner representationWorth et al. (2008), or even hydrodynamic formulations of quantum mechanics à la Bohm-de BroglieBurghardt and Cederbaum (2001a, b); Burghardt and Bagchi (2006). In contrast, the present work adopts a fully classical treatment of the solvent molecules and focuses on clarifying the formally exact framework for a given QM/MM partitioning, as typically addressed in ab initio molecular dynamics.
@@
Despite its considerable achievements, ab-initio molecular dynamics is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space. The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al. (2013); Rozsa et al. (2018). To alleviate ab-initio molecular dynamics computational burden, QM/MMWarshel and Levitt (1976); Mulholland, Lyne, and Karplus (2000); Senn and Thiel (2009a); Groenhof (2013) methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM). The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion. The choice of the classical model to represent molecular entities together with the choice of the QM/MM interaction lacks of a unique definition and consists into a whole field of researchRiniker (2018); Harrison et al. (2018); Unke et al. (2021); Wu et al. (2023); Böselt, Thürlemann, and Riniker (2021); Ho et al. (2025) involving both physically-motivated and machine-learning oriented approaches. However, despite the significant gain in computational cost with respect to ab initio molecular dynamics, sampling the phase space of large systems remains computationally intensive even in QM/MM approaches.
@@
Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al. (2020); Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians. However, just as in molecular dynamics based QM/MM approaches, the coupling interaction between the QM and MM regions lacks of a unique practical definition. The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009b); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023). The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals. While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007) introduced the joint density functional theory which consists in a DFT formulation of the QM/MM problem. The authors obtain a multi component electron/nuclei/solvent DFT formulation, which is justified by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei. However, to derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei, without explicitly indicating how this is done. Therefore, this step makes it difficult to pursue a rigorous derivation. Their working functional is instead defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals. In other workJeanmairet, Levesque, and Borgis (2020); Tang et al. (2020), a similar form of the functional was invoked without further theoretical justification.
@@
Under these assumptions, the primary objective of this paper is to derive, in a straightforward manner, the one-body density functional theory (DFT) formulation for a QM/MM system. Therefore, The key steps of this derivation are as follows: i) We begin by defining the general Hamiltonian of the system, considering a broadly applicable QM/MM Hamiltonian (see Sec. II.1). ii) We then express the exact solution of the problem using full-system quantities (i.e., those describing all particles). This includes the Helmholtz free energy of the QM/MM system, along with the corresponding equilibrium density matrix and partition function (see Sec. II.2). iii) Next, we reformulate the problem as the minimization of a functional over generic full-system density matrices. In our case, this means expressing the Helmholtz free energy as a functional of QM/MM density matrices that describe all particles in the system (see Sec. II.3.1). iv) We then decompose this functional into intrinsic (arising from kinetic energy, inter-particle interactions and entropic contributions) and external (arising from external potentials) contributions (see Sec II.3.2). v) Under certain assumptions about the form of the external potentials, we show how the external part of the functional can be written as a linear functional of one-body densities. In our case, this results in a functional of both classical and quantum one-body densities (see Sec. III.1). vi) Leveraging the many-to-one correspondence between full-system and one-body quantities, we perform a Levy–Lieb constrained search to rewrite the intrinsic free energy functional as a universal functional of the one-body densities (see Sec. III.1). vii) We then establish a connection between electronic DFT (eDFT) and classical DFT (cDFT) by introducing a universal QM/MM correlation functional of the one-body QM and MM densities (see Sec. III.2). viii) Finally, we discuss this formulation in the context of solvation problems and recent developments in QM/cDFT coupling approaches (see Sec. III.3). The Appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ contains a proof of the variational properties of the QM/MM functional introduced here, and the Supplementary Material provides a recall of the Von-Neuman variational principle in statistical QM for the canonical ensemble, an alternative derivation of the QM/MM equilibrium density previously established in the adiabatic basis by Nielsen et. al. in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), and an extension to the semi grand canonical ensemble of the present formalism. As this work is intended for both the QM and MM communities, which often differ in terminology and conceptual frameworks, we highlight certain elements throughout the text to aid accessibility. Some of these clarifications may appear self-evident to some readers; we hope, however, that they do not unduly burden the readability of the article.
@@ II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
We define in Sec. II.1 the working Hamiltonian that describes the QM/MM system. Again, we do not discuss the validity of the partitioning or the quality of the coupling term, we assume that these preliminary considerations have already been addressed. The equilibrium density of a mixed QM/MM system is then given in Sec. II.2, and we briefly discuss the physical meaning of that result. We eventually show in Sec. II.3 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
II.1 The Hamiltonian of the quantum-classical system
@@
In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous. In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons). These particles interact with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$, which is local in the position representation, and with each other through a pairwise potential ${w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j})$. The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
@@
II.2 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy

### v6

@@ Abstract
Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures. However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations , the two major ones being: i) the definition of the QM and MM regions as well as the description of their coupling, and ii) the choice of the methods and levels of approximation made to describe each region.
This paper addresses the second point and develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation. We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble. We start by recalling the expression of the adiabatic equilibrium density matrix for a mixed system made of $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles, which can be related to a partial Wigner transformation. Then, we propose a variational formulation of the Helmholtz free energy in terms of the full , non-equilibrium, QM/MM density matrix. Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities. Therefore, this paper generalizes both cDFT and electronic DFT (eDFT) to QM/MM systems.
@@ I Introduction
Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications. Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion due to the typically large gaps in electronic energies. In this context, because of the large mass of nuclei with respect to that of electrons, it is physically relevant to maintain electrons at the quantum level and to treat nuclei as classical particles. Nevertheless, the forces acting on the nuclei are computed from the solutions of the quantum electronic ground state, i.e. from the solutions of the Born Oppenheimer Hamiltonian. This forms the basis of ab-initio molecular dynamics Iftimie et al. (2005); Silvestrelli and Parrinello (1999) which recovers the classical nuclear entropic contributions through either classical Monte Carlo algorithms or classical trajectories under the ergotic hypothesis.
@@
Nevertheless, it is important to keep in mind that nuclear quantum effects play a significant role for light nuclei and can substantially alter the rates of certain chemical processes, such as proton transfer reactionsBała et al. (1996); Hanna and Kapral (2005); Kim and Hammes-Schiffer (2006); Costentin et al. (2010); Hammes-Schiffer and Stuchebrukhov (2010). Moreover, such proton transfers can also be strongly influenced by the solvation effects, specially when polar solvents are usedKapral (2015). In this context, two fundamentally different yet interconnected nuclear quantum effects are at work: (i) the intrinsic delocalization of nuclear wave packets due to the quantum nature of nuclei, and (ii) the nonadiabatic couplings that arise between various adiabatic electronic states. Nuclear delocalization can be captured through explicit nuclear wave functions (see, for example, Ref. Beck et al., 2000 for a review of time-dependent mean-field approaches, and Ref. Hammes-Schiffer, 2021 for the nuclear-electron orbital approach) or via path-integral formalisms such as ring-polymer molecular dynamicsHabershon et al. (2013). Nonadiabatic couplings, on the other hand, are commonly addressed using surface hopping methods based on classical nuclear trajectoriesTully (1990); Hammes-Schiffer and Tully (1994); Ben-Nun and Martı’nez (2000); Li et al. (2005); Richardson and Thoss (2013); Ananth (2013), modified time-dependent mean-field techniquesWorth et al. (2008), semiclassical expansions of the density matrix and its Wigner representationWorth et al. (2008), or even hydrodynamic formulations of quantum mechanics à la Bohm-de BroglieBurghardt and Cederbaum (2001a, b); Burghardt and Bagchi (2006). In contrast, the present work adopts a fully classical treatment of the solvent molecules and focuses on clarifying the formally exact framework for a given QM/MM partitioning, as typically addressed in ab initio molecular dynamics.
@@
Despite its considerable achievements, ab-initio molecular dynamics is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space. The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al. (2013); Rozsa et al. (2018). To alleviate ab-initio molecular dynamics computational burden, QM/MMWarshel and Levitt (1976); Mulholland et al. (2000); Senn and Thiel (2009); Groenhof (2013) methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM). The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion. The choice of the classical model to represent molecular entities together with the choice of the QM/MM interaction lacks of a unique definition and consists into a whole field of researchRiniker (2018); Harrison et al. (2018); Unke et al. (2021); Wu et al. (2023); Böselt et al. (2021); Ho et al. (2025) involving both physically-motivated and machine-learning oriented approaches. However, despite the significant gain in computational cost with respect to ab initio molecular dynamics, sampling the phase space of large systems remains computationally intensive even in QM/MM approaches.
@@
Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al. (2020); Petrosyan et al. (2005, 2007); Jeanmairet et al. (2020); Labat et al. (2024). In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians. However, just as in molecular dynamics based QM/MM approaches, the coupling interaction between the QM and MM regions lacks of a unique practical definition. The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009); Giovannini et al. (2017); Clemente et al. (2023). The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals. While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
Petrosyan and co-workersPetrosyan et al. (2005, 2007) introduced the joint density functional theory which consists in a DFT formulation of the QM/MM problem. The authors obtain a multi component electron/nuclei/solvent DFT formulation, which is justified by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei. However, to derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei, without explicitly indicating how this is done. Therefore, this step makes it difficult to pursue a rigorous derivation. Their working functional is instead defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals. In other workJeanmairet et al. (2020); Tang et al. (2020), a similar form of the functional was invoked without further theoretical justification.
@@
Under these assumptions, the primary objective of this paper is to derive, in a straightforward manner, the one-body density functional theory (DFT) formulation for a QM/MM system. Therefore, The key steps of this derivation are as follows: i) We begin by a qualitative discussion in Sec. II.1 sketching how the mathematical objects of QM/MM systems can be obtained by Wigner transformations, and then define properly in Sec. II.2 the general form of the Hamiltonian of the system, considering a broadly applicable QM/MM Hamiltonian. ii) We then express the exact solution of the problem using full-system quantities (i.e., those describing all particles). This includes the Helmholtz free energy of the QM/MM system, along with the corresponding equilibrium density matrix and partition function (see Sec. II.3). iii) Next, we reformulate the problem as the minimization of a functional over generic full-system density matrices. In our case, this means expressing the Helmholtz free energy as a functional of QM/MM density matrices that describe all particles in the system (see Sec. II.4.1). iv) We then decompose this functional into intrinsic (arising from kinetic energy, inter-particle interactions and entropic contributions) and external (arising from external potentials) contributions (see Sec II.4.2). v) Under certain assumptions about the form of the external potentials, we show how the external part of the functional can be written as a linear functional of one-body densities. In our case, this results in a functional of both classical and quantum one-body densities (see Sec. III.1). vi) Leveraging the many-to-one correspondence between full-system and one-body quantities, we perform a Levy–Lieb constrained search to rewrite the intrinsic free energy functional as a universal functional of the one-body densities (see Sec. III.1). vii) We then establish a connection between electronic DFT (eDFT) and classical DFT (cDFT) by introducing a universal QM/MM correlation functional of the one-body QM and MM densities (see Sec. III.2). viii) Finally, we discuss this formulation in the context of solvation problems and recent developments in QM/cDFT coupling approaches (see Sec. III.3). The Appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ contains a proof of the variational properties of the QM/MM functional introduced here, and the Supplementary Material provides a recall of the Von-Neuman variational principle in statistical QM for the canonical ensemble, an alternative derivation of the QM/MM equilibrium density previously established in the adiabatic basis by Nielsen et. al. in the context of QM/MM dynamicsNielsen et al. (2001), and an extension to the semi grand canonical ensemble of the present formalism. As this work is intended for both the QM and MM communities, which often differ in terminology and conceptual frameworks, we highlight certain elements throughout the text to aid accessibility. Some of these clarifications may appear self-evident to some readers; we hope, however, that they do not unduly burden the readability of the article.
@@ II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
As the formalism uses mixed QM/MM objects that might seem rather odd in their mathematical design, we begin this derivation by providing in Sec.II.1 a qualitative discussion on how one can obtain these QM/MM objects thanks to the concept of Wigner transform. We then define in Sec. II.2 the working Hamiltonian that describes the QM/MM system. Again, we do not discuss the validity of the partitioning or the quality of the coupling term, we assume that these preliminary considerations have already been addressed. The equilibrium density of a mixed QM/MM system is then given in Sec. II.3, and we briefly discuss the physical meaning of that result. We eventually show in Sec. II.4 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
II.1 Qualitative discussion: dealing with both QM and MM particles thanks to the Wigner transform
Properly defining objects for a QM/MM system can appear as a priori relatively odd as the mathematical spaces of the two theories are very different. We provide here a qualitative discussion (which is fully expanded in the supplementary materials) on how one can obtain naturally QM/MM objects with the help of the so-called Wigner-Weyl transformation.
Observables in classical statistical mechanics are obtained as integrals over the $(Q,P)$ classical phase space,
$\langle O\rangle_{f}=\int dQdPf(Q,P)O(Q,P),$ (1)
where $f(Q,P)$ is a classical probability distribution and $O(Q,P)$ is a classical observable, while the QM theory uses density matrices $\hat{\rho}$ and potentially non local operators $\hat{O}$ acting in $\mathbb{R}\times\mathbb{R}$, namely
$\langle\hat{H}\rangle_{\hat{\rho}}=\int dxdx^{\prime}\hat{\rho}(x,x^{\prime})\hat{O}(x^{\prime},x),$ (2)
where $\hat{\rho}(x,x^{\prime})$ and $\hat{O}(x^{\prime},x)$ are the real-space representation of the density matrix $\hat{\rho}$ and $\hat{O}$ operators, respectively. Nevertheless, one can notice that the $(Q,P)$ classical phase space is isomorph to the $(x,x^{\prime})$ representations of the QM operators. The Wigner-Weyl transformWigner (1932) exploits such a similarity and allows then for a phase space representation of QM density matrices and operators (see Ref. Case, 2008 for a pedagogical introduction).
More precisely, for any QM operator $\hat{O}$ one can define its Wigner transformation labelled $\tilde{O}_{W}$, which is now a function of the $(Q,P)$ variables. One can also define the Wigner transformed density matrix $\tilde{\rho}_{W}(Q,P)$ associated to $\hat{\rho}$, which allows then to write the QM observable of Eq. (2) as a classical phase space integral, namely
$\langle\hat{O}\rangle_{\hat{\rho}}=\int dQdP\tilde{\rho}_{W}(Q,P)\tilde{O}_{W}(Q,P).$ (3)
An important property of this framework is that QM operators expressed as ”simple” functions of the QM operators $\hat{Q}$ and $\hat{P}$ (i.e. functions with no products of $\hat{P}$ and $\hat{Q}$) are transformed as the corresponding purely classical function. An important special case of this property is the Hamiltonian operator whose Wigner transformation is precisely the classical Hamiltonian, i.e.
$\tilde{H}_{W}(Q,P)=\frac{P^{2}}{2M}+V(Q)\equiv H(Q,P).$ (4)
Therefore, the expectation value of the QM Hamiltonian over a QM density matrix is obtained as
$\langle\hat{H}\rangle_{\hat{\rho}}=\int dXdP\tilde{\rho}_{W}(Q,P)H(Q,P),$ (5)
which is very similar to the definition of the statistical classical mechanics. Nevertheless the quantum nature of the problem remains encoded in the fact that the object $\tilde{\rho}_{W}(Q,P)$ is potentially complex-valued in opposition to classical probability distributions.
[…]
In the case now where two types of particles are present in the system with a clear difference of masses, one can perform a Wigner transformation only on the heavy particles and one obtains a power series in terms of the ratio of the masses $\alpha\ll 1$. As shown originally by Nielsen et. al.Nielsen et al. (2001), the zeroth-order term in $\alpha$ yields to a classical description of the heavy particles, while maintaining the full QM description of the light particles. The obtained semi-Wigner transformed density matrix corresponds then to a QM/MM description of the system: the heavy particles are described by the classical variables $(Q,P)$ while the light particles are described by a non local operator. An important aspect of these results is that the QM operators depend parametrically on the $(Q,P)$ classical variables. We will use this important result as the starting of our derivation (see Sec. II.3) which will allow us to understand how generic QM/MM density matrices can be built in order to design our variational principle.
II.2 The Hamiltonian of the quantum-classical system
@@
Notice that the form of the QM/MM Hamiltonian as written in Eq. (6) can actually be obtained as the partial Wigner transformation of a purely QM Hamiltonian, and where the Wigner transformation is done only on the $N_{{\text{mm}}}$ particles. In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous. In Eq.(6), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons). These particles interact with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$, which is local in the position representation, and with each other through a pairwise potential ${w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j})$. The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
@@
II.3 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy